# Towards Autonomous Driving with Small-Scale Cars: A Survey of Recent Development

### v1

@@ II-A Hardware platforms
The significance of a small-scale smart city platform, accommodating multiple vehicles, is underscored as it plays a crucial role in achieving harmony and efficient collaboration of a fully autonomous driving world. To this end, the University of Delaware Smart Scaled City (UDSSC) [16], developed for education and research on connected and autonomous driving, emerges as a distinctive contribution. It features a downscaled city environment with diverse traffic scenarios, including intersections and roundabouts, and incorporates Micro Connected and Automated Vehicles (MCAV). Built on the Pololu Zumo [^2] https://www.pololu.com/category/170/zumo-32u4-oled-robot/ platform, the MCAVs have a differential drive DC motor setup with encoders, an IMU, and an Arduino [^3] https://www.arduino.cc/. Each vehicle is additionally equipped with a Raspberry Pi as its compute unit, a line-following sensor, and a front-facing distance sensor. Notably, the absence of a camera or LiDAR prompts the utilization of a central localization system for accurate positioning. UDSSC serves a specialized purpose, exclusively engaging in cooperative and connected driving tasks such as merging [16] and roundabout traversal [57, 58, 59]. The platform is unfortunately not accompanied by any open-source code or hardware accessibility and is not available for purchase.
For the sake of brevity, we have condensed the details of the usually used platforms into Table I. This table offers a comparative analysis, outlining key characteristics such as steering dynamics, installed sensors, utilized software frameworks, and, when relevant, the simulation software integrated into each platform.
TABLE I: Small-scale car platforms: an overview of the hardware and software setup.
@@
Platforms Size Vehicle Dynamics Sensors Computation Unit Software Framework Simulation Platform Open-source11footnotemark: 1 Commercial
@@
AutoRally [9] 1/5th Ackermann Camera, IMU, GPS, Hall-effect sensor Intel i7-6700 Nvidia GTX-750ti SC ROS Gazebo SW No22footnotemark: 2
ART/ATK [13] 1/6th Ackermann Camera,3D LiDAR Jetson Xavier NX ROS2 Chrono No
BARC [8] 1/10th Ackermann Camera, LiDAR, IMU, GPS ODROID-XU4 ROS No
Donkeycar [11] 1/10th 1/16th Ackermann Camera, LiDAR, IMU, Encoder RPi/Jetson Nano Python Donkeycar Gym SW, HW No22footnotemark: 2
F1TENTH [12] 1/10th Ackermann Camera, LiDAR, IMU Jetson TX2 ROS Gazebo F1TENTH Gym SW No22footnotemark: 2
RACECAR(MIT) [18] 1/10th Ackermann Camera, LiDAR, IMU, Encoder Jetson Tegra X1 ROS Gazebo SW, HW No22footnotemark: 2
MuSHR [19] 1/10th Ackermann Camera, LiDAR, IMU, Bump sensor Jetson Nano ROS SW No22footnotemark: 2
@@
Autominy [23] 1/10th Ackermann Camera, LiDAR, IMU, Encoder Intel NUC ROS Gazebo SW No
@@
Autonomouscar [21] 1/10th Ackermann Camera, LiDAR, IMU, Encoder, ToF Sensor, Indoor GPS RPi 4 ROS, LabVIEW SW, HW No
CoRoLa [24] 1/10th Ackermann Camera, Encoder, Ultrasonic sensor RPi 4 ROS2 Based on LGSVL No
AutoDRIVE [60] 1/14th Ackermann Camera, LiDAR, IMU, Encoder, Indoor GPS Jetson Nano ROS AutoDRIVE Simulator SW, HW No
@@
ITS car [62] 1/18th Ackermann Camera RPi 3 B+ No
µcar [63] 1/18th Ackermann IMU, Encoder RPi Zero W SW, HW No
UDSSC MCAV [16] 1/25th Differential IMU, line following, IR sensor RPi 3 ROS SUMO No
Chronos [64] 1/28th Ackermann IMU, Encoder Espressif ESP32 ROS SW, HW No
Go-CHART [65] 1/28th Differential Camera, LiDAR, Bump sensor RPi 3 ROS SW, HW No
Cambridge Minicar [66] 75x81x197mm Ackermann Indoor positioning system33footnotemark: 3 RPi Zero SW, HW No
ORCA Racer [15] 1/43th Ackermann IMU, Indoor positioning system33footnotemark: 3 ARM Cortex M4 IMU No
@@
GRITSBot [67] 31x30mm Differential IMU, IR sensor Atmega 328 SW, HW No
HydraOne [68] 27x32cm Omni Camera, 3D LiDAR, Encoders Jetson TX2 ROS HW No
Pheeno [31] 10cm ⌀ Differential Camera, IMU, Encoder, IR sensor ATmega328P ARM Cortex-A7 Python No
@@
MarXbot [33] 17cm ⌀ Differential Camera, IMU, IR sensor, 2D force sensor ARM 11 processor ROS Gazebo No
WolfBot [69] 17.5cm ⌀ Omni Camera, IMU, IR sensor, Microphone BeagleBone ROS MATLAB SW, HW No
LabRAT [70] - Differential IR sensor Atmega324p C Player/Stage SW, HW No
mAGV [71] - Differential IMU, Encoder, RFID RPi Zero W ROS No
Rover [72] - Differential Camera, LiDAR, IMU, GPS, Ultrasonic sensor Jetson TK1 ROS Gazebo No
@@
    For open-source types, SW: Software and HW: Hardware.
@@
  • 4.
    Educational robot kit.
  • 5.
    Platform-specified block and visual programming languages.
@@ II-B Simulators
First, for the requirements of simulating the vehicle dynamics and control systems, CarMaker [75] and CarSim [76] emerge as notable choices. They furnish simulation environments capable of comprehensively assessing various facets of vehicle behavior, encompassing handling, braking, and acceleration, across diverse driving conditions. For autonomous driving systems, the sensor models are as important as the physics engine, therefore, Gazebo [77] plays an important role in robotics and related research areas. It provides a modular system that facilitates the integration of diverse sensors and physics models. For small-scale cars such as Turtlebot, Duckietown, DeepRacer, F1TENTH, MuSHR, Autorally, and others, simulations based on Gazebo have been developed. For a macro-scale simulation, SUMO [78] emerges as a platform for in-depth analysis and optimization of transportation systems, road networks, and traffic management strategies. Similarly, Highway-env [79] provides a gym-based environment specifically designed for simulating highway-based road topologies. In the realm of more detailed autonomous driving setups, various open-source simulators have gained popularity, with TORCS [80], CARLA [81], AIRSIM [82], DeepDrive [83], and LGSVL [84] being notable examples. These simulators not only boast high-fidelity physics engines for simulating vehicle dynamics but also support a diverse array of sensors. This allows for hyper-realistic simulations of traffic environments, facilitating rigorous testing of AVs. For multi-agent systems, MADRaS [85], built on top of TORCS, and Flow [86], built on top of SUMO, are commonly used.
Besides the universal simulators, some simulators are specially tailored for small-scale cars. For example, Gym-Duckietown [87] for Duckietown, provides a fast, open, and highly customizable simulator, to train or test various algorithms for different driving tasks. The DeepRacer cloud simulator from Amazon for the DeepRacer racing platform provides an online platform to customize the algorithm, choose the tasks, and modify the structure of the networks. Donkey-Gym [88] for Donkeycar is similar to Gym-Duckietown, but built on the Unity game platform with internal physics and graphics, also the F1TENTH Gym environment [89] for the F1TENTH platform. With these specially designed simulation platforms, the real-world deployment of autonomous small-scale cars will be facilitated after training and testing in the simulation. Nevertheless, the Simulation to Reality (Sim2Real) gap remains a challenge that needs to be addressed before successfully deploying the trained algorithms in the real world.
@@ II-C Sim2Real transfer
Sim2Real is a concept in robotics and AVs that involves transferring skills, knowledge, or models acquired in a simulated environment to real-world applications. Here, the real-world setting or scenario in which the robot is intended to operate and execute tasks is termed the target domain. Conversely, the data, and experiences, shaping the development of the robotic system are called the source domain. The core objective of Sim2Real is to devise algorithms and methodologies capable of effectively bridging the disparities between these two domains, known as the Sim2Real gap. For instance, this gap may manifest as dynamic differences or discrepancies in the sensing part, where the simulated images and the real images are different. Although the Sim2Real transfer is primarily centered on transferring Deep Reinforcement Learning (DRL) policies from simulation to the real world, it can also be more broadly considered as Machine Learning (ML) problems for the sensing part, for the agent facing situations in the real world that have not appeared in simulation. To address the Sim2Real gap, an array of methods is proposed, including system modeling, dynamics randomization, and randomization for sensing. Here we introduce the most used methods for small-scale cars.

### v2

@@ II-A5 Multi Vehicle Platforms
The significance of a small-scale smart city platform, accommodating multiple vehicles, is underscored as it plays a crucial role in achieving harmony and efficient collaboration of a fully autonomous driving world. To this end, the University of Delaware Smart Scaled City (UDSSC) [21], developed for education and research on connected and autonomous driving, emerges as a distinctive contribution. UDSSC features a scaled-down urban environment with diverse traffic scenarios, including intersections and roundabouts, and uses Micro Connected and Automated Vehicles (MCAVs). These vehicles are built on the Pololu Zumo platform and feature a differential-drive DC motor setup with encoders, an IMU, and an Arduino microcontroller. Each MCAV is further equipped with a Raspberry Pi as its compute unit, a line-following sensor, and a front-facing distance sensor. Notably, the absence of onboard cameras or LiDAR necessitates the use of a centralized localization system for precise positioning. The platform is tailored for cooperative and connected driving tasks, such as merging [21] and roundabout navigation [56, 57, 58]. However, it is constrained by several limitations. The reliance on an external localization system ties the platform to a specific physical testbed, as relocating the system requires substantial effort. Additionally, the lack of onboard sensors for localization restricts its use to studies on multi-vehicle interactions. Unfortunately, the UDSSC platform is not open source; its hardware and software are proprietary, and it is not available for purchase. Moreover, the absence of detailed documentation limits its usability. As a result, the platform primarily serves as inspiration for developing custom solutions with similar capabilities.
For the sake of brevity, the details of commonly used platforms have been summarized in tables. Table II presents a comparative hardware analysis of various platforms, highlighting key characteristics such as steering dynamics, actuators, installed sensors, and prices, whether for crafting or purchasing a platform. Table III focuses on the software aspect, detailing the software frameworks employed and, where applicable, the simulation tools integrated into each platform. Additionally, the primary focus tasks of each platform are provided as a reference for those interested in exploring specific driving behaviors.
TABLE II: Small-scale car platforms: an overview of the hardware setup.
@@
Platforms Size Vehicle Dynamics Actuator Sensors Computation Unit Runtime Commercial available Price (USD)
@@
AutoRally [11] 1/5th Ackermann Two servo motors Camera, IMU, GPS, Hall-effect sensor Intel i7-6700 Nvidia GTX-750ti SC <1h
@@
2613
77footnotemark: 7
10k
ART/ATK [18] 1/6th Ackermann One brushless DC motor One servo motor Camera, 3D LiDAR Jetson Xavier NX
@@
2613
BARC [13] 1/10th Ackermann One brushless DC motor Camera, LiDAR, IMU, GPS ODROID-XU4
[…]
2613
@@
2613
77footnotemark: 7
350
F1TENTH [12] 1/10th Ackermann One brushless DC motor Camera, LiDAR, IMU Jetson TX2 <1h
@@
2613
77footnotemark: 7
3800
RACECAR(MIT) [24] 1/10th Ackermann One brushless DC motor One servo motor Camera, LiDAR, IMU, Encoder Jetson Tegra X1
@@
2613
77footnotemark: 7
2600
MuSHR [25] 1/10th Ackermann One brushless DC motor One servo motor Camera, LiDAR, IMU, Bump sensor Jetson Nano
@@
Autominy [27] 1/10th Ackermann One brushless DC Servomotor Camera, LiDAR, IMU, Encoder Intel NUC
@@
2613
[…]
Autonomouscar [26] 1/10th Ackermann One brushed DC motor Camera, LiDAR, IMU, Encoder, ToF Sensor, Indoor GPS RPi 4
[…]
2613
CoRoLa [28] 1/10th Ackermann One brushless DC motor One servo motor Camera, Encoder, Ultrasonic sensor RPi 4
[…]
2613
AutoDRIVE [59] 1/14th Ackermann Two DC geared motors Camera, LiDAR, IMU, Encoder, Indoor GPS Jetson Nano
[…]
2613
@@
2613
UDSSC MCAV [21] 1/25th Ackermann One geared DC motor IMU, line following, IR sensor RPi 3 90m
[…]
2613
[…]
2613
[…]
2613
Cambridge Minicar [63] 75x81x197mm Ackermann Indoor positioning system88footnotemark: 8 RPi Zero 2h
[…]
2613
ORCA Racer [20] 1/43th Ackermann IMU, Indoor positioning system88footnotemark: 8 ARM Cortex M4 20m
[…]
2613
@@
[…]
2613
HydraOne [65] 27x32cm Omni Two encoder motors Camera, 3D LiDAR, Encoders Jetson TX2
[…]
2613
77footnotemark: 7
7200
Pheeno [31] 10cm ⌀ Differential Two micro gear motors Camera, IMU, Encoder, IR sensor ATmega328P ARM Cortex-A7 5h
[…]
2613
77footnotemark: 7
270
[…]
MarXbot [32] 17cm ⌀ Differential Camera, IMU, IR sensor, 2D force sensor ARM 11 processor
[…]
2613
WolfBot [66] 17.5cm ⌀ Omni Camera, IMU, IR sensor, Microphone BeagleBone
[…]
2613
LabRAT [67] - Differential Two DC gearmotors IR sensor Atmega324p 3h
[…]
2613
[…]
@@
    http://donkeycar.com
@@
    https://www.quanser.com/products/qcar
  • 3.
    https://github.com/NVIDIA-AI-IOT/jetracer
  • 4.
    https://www.waveshare.com/wiki/PiRacer_AI_Kit
  • 5.
    https://www.turtlebot.com
  • 6.
    https://jetbot.org
  • 7.
@@
[…]
2 https://github.com/tawnkramer/gym-donkeycar.
3 https://github.com/duckietown/gym-duckietown.
4 https://github.com/f1tenth/f1tenth_gym.
TABLE III: Software setups and target tasks for different small-scale car platforms.
@@ II-B Simulators
First, for the requirements of simulating the vehicle dynamics and control systems, CarMaker and CarSim are widely recognized. These platforms provide environments for assessing various aspects of vehicle behavior, including handling, braking, and acceleration, under diverse driving conditions. For autonomous systems, sensor models are as crucial as physics engines. Gazebo [68] is a prominent open-source platform in robotics, offering modular systems for integrating diverse sensors and physics models. Gazebo-based simulations have been developed for small-scale platforms like Turtlebot, Duckietown, DeepRacer, F1TENTH, MuSHR, and Autorally, demonstrating its flexibility and adaptability to different research needs. Other robotic simulators such as ARGoS and V-REP are also used as bases for small-scale car platforms [17, 22]. For macro-scale simulations, SUMO [69] is a leading platform for analyzing transportation systems, optimizing road networks, and developing traffic management strategies. It has been used for training traffic control algorithms in multi-vehicle systems, as discussed in [58]. In the realm of more detailed autonomous driving setups, various open-source simulators have gained popularity, with TORCS [70], CARLA [71], AIRSIM [72], and LGSVL [73] being notable examples. These platforms feature high-fidelity physics engines and extensive sensor support, enabling hyper-realistic simulations of traffic environments. They are particularly effective for training larger small-scale vehicles, such as 1/5th scale models, before real-world deployment [74, 75, 76, 77]. In addition to general-purpose simulators, several platforms are tailored to specific small-scale systems. For example, Gym-Duckietown provides a fast and customizable environment for Duckietown, supporting various driving tasks. The DeepRacer cloud simulator by Amazon offers an online platform for customizing algorithms and testing them within the DeepRacer ecosystem. Donkey-Gym, designed for Donkeycar, uses the Unity game engine for enhanced physics and graphics. F1TENTH Gym is optimized for F1TENTH platforms, offering specialized simulation capabilities. AutoDRIVE simulator provides a comprehensive environment for various driving tasks for AutoDRIVE cars. For multi-agent systems, platforms like MADRaS [78] (built on TORCS) and Flow [79] (built on SUMO) are frequently employed. These tools facilitate the study of interactions and behaviors in complex traffic and multi-agent environments. In Table III, a summary of simulators used for main small-scale car platforms is shown.
Communication in robotics is as critical as the simulator itself, and ROS [80] serves as a key middleware platform for managing robotic software development. ROS facilitates communication between various robotic components, including sensors, actuators, and decision-making algorithms, enabling seamless integration and scalability. It also offers a vast collection of pre-built packages and libraries for common tasks like motion planning and image processing. Currently, ROS exists in two versions: ROS 1 and ROS 2. ROS 1, the original version, is widely used and supported by the community; however, its centralized architecture limits real-time capabilities and lacks robust security features. In contrast, ROS 2 addresses these limitations with a decentralized design based on the Data Distribution Service, offering enhanced scalability, real-time support, and Quality of Service. While ROS 1 remains suitable for projects reliant on mature tools and is ideal for educational or non-real-time applications, ROS 2 is recommended for new projects requiring real-time performance, high scalability, security, and cross-platform compatibility. Nevertheless, the Simulation to Reality (Sim2Real) gap remains a challenge that needs to be addressed before successfully deploying the trained algorithms in the real world.
@@ II-C Sim2Real Transfer
Sim2Real is a concept in robotics and AVs that involves transferring skills, knowledge, or models acquired in a simulated environment to real-world applications. Here, the real-world setting or scenario in which the robot is intended to operate and execute tasks is termed the target domain. Conversely, the data, and experiences, shaping the development of the robotic system are called the source domain. The core objective of Sim2Real is to devise algorithms and methodologies capable of effectively bridging the disparities between these two domains, known as the Sim2Real gap. For instance, this gap may manifest as dynamic differences or discrepancies in the sensing part, where the simulated images and the real images are different. Although the Sim2Real transfer is primarily centered on transferring Deep Reinforcement Learning (DRL) policies from simulation to the real world, it can also be more broadly considered as ML problems for the sensing part, for the agent facing situations in the real world that have not appeared in simulation. To address the Sim2Real gap, an array of methods is proposed, including system modeling, dynamics randomization, and randomization for sensing. Here we introduce the most used methods for small-scale cars.